OpenStudy (anonymous):

a toy rocket is fired from the ground with an upward velocity of 200 feet per second. How many seconds will the rocket take to return to the ground?

OpenStudy (freckles):

can you give the equation for the position of the rocket

OpenStudy (anonymous):

h=-16t^2 +v0t+h0, where h is the final height in feet, t is the time of travel in seconds,v0 is the initial velocity in feet per second,and h0 is the initial height in feet of an object traveling upward

OpenStudy (freckles):

stupid question upward velocity means not initial right? and if not, what is the relationship between upward and initial velocity? or what does upward velocity mean exactly?

OpenStudy (freckles):

@ganeshie8 can you tell me what upward velocity means exactly?

OpenStudy (anonymous):

it does not specify all i am given is the question and the formula

OpenStudy (freckles):

and pretty please :)

ganeshie8 (ganeshie8):

yes its same as the initial velocity before leaving the ground

OpenStudy (freckles):

oh that makes this easier for me to read thank you so much

OpenStudy (freckles):

well we know the rocket started with initial height 0 so we know h0 is 0

OpenStudy (freckles):

\[h=-16t^2+200t\]

OpenStudy (freckles):

we want to know when h is zero again

OpenStudy (freckles):

besides at t=0

OpenStudy (freckles):

0=-16t^2+200t solve for t

OpenStudy (anonymous):

thank you and one more question if you could?

OpenStudy (freckles):

post

OpenStudy (anonymous):

(a-1)(a+4)=14 p.s I'm really bad at math:(

OpenStudy (freckles):

you will have to make it ugly before you make it pretty

OpenStudy (freckles):

you need everything on one side

OpenStudy (freckles):

and multiply the (a-1)(a+4) and then combine like terms

OpenStudy (freckles):

\[(a-1)(a+4)=14 \\ \text{ subtract 14 on both sides } \\ (a-1)(a+4)-14=0 \\ \text{ multiply the } (a-1)(1+4) \text{ then subtract the 14 there }\]

OpenStudy (anonymous):

thank you so much :)

OpenStudy (freckles):

np :)

OpenStudy (freckles):

like if you still need help on solving it let me know

OpenStudy (freckles):

i will be here for a few more minutes

OpenStudy (anonymous):

I'm done for tonight but thanks for all your help

OpenStudy (freckles):

np :) gave a good night

OpenStudy (freckles):

have*

OpenStudy (anonymous):

thanks, if you need help in english i can help lol thats my best subject

OpenStudy (freckles):

i feel like you are making fun of my english a little bit :p just kidding

OpenStudy (anonymous):

you can make fun of my math :)

OpenStudy (freckles):

i might

OpenStudy (anonymous):

:) night